OpenStudy (anonymous):

What cosine function represents an amplitude of 2, a period of 2π, a horizontal shift of π, and a vertical shift of −1?

OpenStudy (anonymous):

f(x) = −1 cos πx + 2 f(x) = −1 cos (x − π) + 2 f(x) = 2 cos (x − π) − 1 f(x) = 2 cos πx − 1

OpenStudy (anonymous):

@iGreen @midhun.madhu1987 @hartnn

OpenStudy (phi):

can you rule out any of the choices?

OpenStudy (anonymous):

A and B?

OpenStudy (phi):

A and B are out. they show an amplitude of 1 (we don't care about the sign of -1)

OpenStudy (anonymous):

alright, so whats next?

OpenStudy (phi):

If you can remember this if you have cos( A x) where A is a number and you want to find the period of cos(Ax) match up \[ A= \frac{2 \pi}{T} \] and solve for T (the period) in your case, they say a period of 2π, i.e. T= 2π so solve \[ A = \frac{2 \pi}{2 \pi} \] for A

OpenStudy (anonymous):

So how do we get the answer? @phi

OpenStudy (anonymous):

Is it D? @phi

OpenStudy (agent0smith):

Horizontal shift means you'd need to add or subtract pi in the brackets of cos( x + ? )

OpenStudy (anonymous):

is it d? @agent0smith

OpenStudy (agent0smith):

See last post